Grade 11General Physics

A river is flownig west to east at a speed of 5m/s. A man on the south bank of the river capable of swimming at 10 m/s in still water wants to cross the river without drifting. At what angle with the vertical should he swim to do so?

Rituraj Singh
12 Years agoGrade 11

1 Answer

Arun Kumar
12 Years ago
Hello Student,
Well for that net horizontal velocity =0
=>vcos\theta=5
=>cos\theta=0.5
=>\theta=30
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty